# Generating Shrinkers by Mean Curvature Flow

### v2

@@ Abstract.
F. Martín’s research is partially supported by the framework of IMAG-María de Maeztu grant CEX2020-001105-M funded by MCIN/AEI/ 10.13039/50110001103. Some of this work was carried out while the authors were visitors at the Simons Laufer Mathematical Sciences Institute (SLMath, formerly MSRI) in Berkeley, CA during the Fall 2024 semester, in a program supported by National Science Foundation Grant No. DMS-1928930.
@@ 1. Introduction
Figure 1. A cross section of a numerical simulation of one of the shrinkers constructed in Section 7. Figure courtesy of Mario Schulz.
@@ 2. Preliminaries
In this paper, “mean curvature flow” means “unit-regular Brakke flow”. Roughly speaking, a unit-regular Brakke flow is an integral Brakke flow that is smooth (without sudden vanishing) in a spacetime neighborhood of each point of Gauss density $1$. (See [white-boundary]*Definition 11.1 for the precise definition.) Any weak limit of unit-regular flows is also unit-regular [white-local].
If $K$ is a closed subset of $\mathbf{R}^{3}$, or, more generally, of a smooth Riemannian $3$-manifold $N$, we let $K(t)$ denote the result of flowing $K$ for time $t$ by the level set flow.
Let $\operatorname{T_{\textnormal{fat}}}:=\operatorname{T_{\textnormal{fat}}}(K)$ be the fattening time, i.e, the infimum of times $t$ such that $K(t)$ has nonempty interior.
Definition 1.
If $M$ is smooth, a regular time for $M$ is a time $t\in[0,\operatorname{T_{\textnormal{fat}}}(M)]$ such that $M(\tau)$ is smooth for all $\tau$ in some neighborhood $[0,\operatorname{T_{\textnormal{fat}}}]\cap(t-\epsilon,t+\epsilon)$ of $t$.
If $M$ is smoothly embedded and compact, then 00 is a regular time. By [white-genus], if $t_{1}<t_{2}<\operatorname{T_{\textnormal{fat}}}(M)$ are regular times, then
@@
By [bamler-kleiner], almost every time in $[0,\operatorname{T_{\textnormal{fat}}}(M)]$ is regular, and all tangent flows at times $t\in(0,\operatorname{T_{\textnormal{fat}}}]$ are given by smoothly embedded shrinkers of multiplicity $1$. Thus if $p\in N$ and $T\in(0,\operatorname{T_{\textnormal{fat}}}(M)]$, then every sequence $t_{i}\uparrow T$ has a subsequence such that
@@
converges smoothly and with multiplicity $1$ on compact subsets to a shrinker $\Sigma$. If $\tau_{i}/t_{i}$ is bounded away from 00 and away from $\infty$, then
@@
We expect there to be only one shrinker at each spacetime point, but that is not known in all cases. (See [chodosh-schulze] for cases in which uniqueness is known.) However, all the shrinkers at $(p,T)$ have the same genus (by Theorem 64 in the appendix), so the following definition makes sense.
@@
At a singular point of genus 00, the shrinker is either a sphere or a cylinder [brendle]. (Earlier, Kleene and Møller proved the result assuming an axis of rotational symmetry [kleene-moller].) Furthermore, the shrinker is unique, by [huisken] (in the spherical case) and [colding-m] (in the cylindrical case). See [zhu] for a simpler proof in the cylindrical case.
@@
If $\operatorname{T_{\textnormal{fat}}}<\infty$, then there is a spacetime singular point $(p,\operatorname{T_{\textnormal{fat}}})$ of positive genus.
@@
We let $T_{\textnormal{pos}}(M)$ be the infimum of times $t$ at which there is a singularity of positive genus.
The set of singular spacetime points of genus $>0$ is a closed subset of spacetime. Thus if $T_{\textnormal{pos}}(M)<\infty$, then there is spacetime singularity $(p,T_{\textnormal{pos}}(M))$ of positive genus. (In [ccms]*§11, $T_{\textnormal{pos}}(M)$ is written $T_{\textnormal{gen}}(M)$.)
@@
Suppose $t\in[0,T)\mapsto M(t)$ is a mean curvature flow in an open set $U$ such that $M(0)$ is a smooth, propertly embedded, multiplicity $1$ submanifold of $U$. Let $K$ be a compact region in $U$ with smooth boundary transverse to $M(0)$. There is an $\epsilon>0$ with the following properties.
@@
$K^{\prime}$ =$=\{\operatorname{dist}(\cdot,K)\leq\delta\},$
@@
Suppose $p_{i}$ converges to $p\in M(0)$ and that $t_{i}>0$ converges to 00. Let $q_{i}$ be the point in $M(0)$ closest to $p_{i}$. Let
@@
(*) The flow $t\in[0,2\epsilon]\mapsto M(t)\cap K^{\prime}$ is smooth.
Assertion (1) follows immediately. Assertion (2) follows from (*2) and the Local Regularity Theorem (Theorem 7).
@@ 3. Oriented Flows
Suppose that $t\in I\mapsto M(t)$ is a unit-regular Brakke flow. We let $\operatorname{Reg}(t)$ be the set of points $p$ such that the flow is smooth and with multiplicity $1$ in a spacetime neighborhood of $(p,t)$. Equivalently, $\operatorname{Reg}(t)$ is the set of points $p$ such that the Gauss Density at $(p,t)$ is $1$. We let $\operatorname{Sing}(t)$ be the set of $p$ such that the Gauss Density of at $(p,t)$ is $>1$.
@@ 4. Platonic and Regular Polygon Shrinkers
    $\Sigma$ occurs as a shrinker in a mean curvature flow $M(\cdot)$ for which $M(0)$ is a compact, smoothly embedded surface in $Q^{c}$.
@@
If $M\in\mathscr{A}$ and if $t<\min\{T_{Q},\operatorname{T_{\textnormal{fat}}}\}$ is a regular time of the flow $M(\cdot)$, then $M(t)\in\mathscr{A}$.
@@
The theorem is a trivial consequence of the facts that (1) level set flow preserves symmetries, and (2) for any compact smooth embedded surface, $\operatorname{genus}(M(t))$ is a decreasing function of regular times in $[0,\operatorname{T_{\textnormal{fat}}}]$. (See (1).) ∎
@@
If $M\in\mathscr{A}$, then there is a connected component $\Sigma$ of $M$ such that the convex hull of $\Sigma$ contains points of $Q$,
@@
Let $M$ be a $G$-invariant surface properly embedded in $\mathbf{R}^{3}$, disjoint from $Q$, and of genus $\leq g$. Suppose $M$ has a connected component $\Sigma$ whose convex hull contains points of $Q$. Then $\Sigma$ has genus $g$, and if $\Omega$ is a simply connected open subset of $Q^{c}$, then $\operatorname{genus}(M\cap\Omega)=0$.
Proof.
@@
$\Psi$ contains points in $F_{12}$ and points in $F_{13}$
@@
Suppose not. Then (by relabelling) we may assume that it contains no points in the face $F_{12}$. It follows, by iterated reflections in the planes in $\mathscr{P}$ containing $E_{3}$, that $\Sigma$ lies in $\operatorname{interior}(K_{3})$, a convex open set disjoint from $Q$. But that contradicts the hypothesis on $\Sigma$. Thus the claim is proved.
Let $M^{\prime}$ be the union of $\Sigma$ and its images under $G$. Let $U^{\prime}$ be the region bounded by $M^{\prime}$ that is disjoint from $Q$. (Thus, a point in $(M^{\prime})^{c}$ is in $U^{\prime}$ if and only if there is a path from that point to a point in $Q$ that is transverse to $M^{\prime}$ and that intersects $M^{\prime}$ in an odd number of points.) Let $K=M^{\prime}\cup U^{\prime}$. Thus $K$ is smooth $3$-manifold-with-boundary, the boundary being $M^{\prime}$.
@@
Let $\gamma$ be an embedded curve in $\Psi$ such that one of its endpoints is in $F_{12}$ and the other endpoint is in $F_{13}$. Iterated reflections in the planes of $\mathscr{P}$ containing $E_{1}$ produces a simple closed curve in $M^{\prime}\cap K_{1}$ that winds once around the line containing $E_{1}$. Of course, there is such a curve for each ray of $Q$. Let $\mathscr{G}$ be the set of those $g+1$ curves. Note that any $g$ of those $g+1$ curves form a basis for the vector space $H_{1}(Q^{c};\mathbf{Z}_{2})$.
Thus if $\iota$ is the inclusion of $K$ into $Q^{c}$, then the map
(2) $\iota_{\#}:H_{1}(K;\mathbf{Z}_{2})\to H_{1}(Q^{c};\mathbf{Z}_{2}).$
is surjective.
Now
$g\geq\operatorname{genus}(M^{\prime})=\dim H_{1}(K;\mathbf{Z}_{2})\geq\dim H_{1}(Q^{c};\mathbf{Z}_{2})=g.$
Thus all the inequalities are equalities and the map (2) is an isomorphism.
Now the $g+1$ curves of $\mathscr{G}$ taken together are homologically trivial (mod $2$) in $Q^{c}$ and thus also in $K$. Thus they bound a surface in $K$. That surface must be connected since no subcollection of those curves bounds a surface in $Q^{c}$. Hence $K$ is connected. Therefore $M^{\prime}=\partial K$ is connected, so $\Sigma=M^{\prime}$.
@@
It remains to show that if $W$ is a simply connected open subset of $Q^{c}$, then $M\cap W$ has genus 00. Since $M$ and $\Sigma$ have genus $g$, $M\setminus\Sigma$ has genus 00, and therefore $(M\setminus\Sigma)\cap W$ has genus 00. Thus it suffices to prove that $\Sigma\cap W$ has genus 00. Suppose to the contrary that $\Sigma\cap W$ has positive genus. Then, by elementary topology (see Lemma 16 below), $\Sigma\cap W$ contains a simple closed curve that does not bound a surface in $K$. Therefore it does not bound a surface in $Q^{c}$ since the map (2) is an isomorphism, and therefore it does not bound a surface in $W$. But that contradicts the assumption that $W$ is simply connected. ∎
@@
Let $K$ be a closed region in $\mathbf{R}^{3}$ with smooth boundary $\Sigma$. Let $W$ be an open subset of $\mathbf{R}^{3}$ such that $\Sigma\cap W$ has positive genus. Then $\Sigma\cap W$ contains a simple closed curve that is homologically nontrivial in $K$.
@@
Let $\alpha_{1}$ and $\alpha_{2}$ be a pair of simple closed curves in $\Sigma\cap W$ that intersect transversely and at exactly one point. We claim that at most one of $\alpha_{1}$ and $\alpha_{2}$ can be homologically trivial in $K$. For otherwise, there would be smooth, compact embedded surfaces $\sigma_{i}$ in $K$ such that $\partial\sigma_{i}=\alpha_{i}$. We can choose the $\sigma_{i}$ so that they intersect transversely and so that they are never tangent to $\Sigma$. Thus $\sigma_{1}\cap\sigma_{2}$ is a finite collection of curves, and so they have an even number of endpoints. But the set of their endpoints is $\alpha_{1}\cap\alpha_{2}$, which has only one point, a contradiction. ∎
@@
Suppose $M\in\mathscr{A}$. Then $T_{Q}\leq T_{\textnormal{pos}}$.
@@
Let $0<T<\min\{T_{Q},T_{\textnormal{pos}}\}$. Let $(p,T)$ be a spacetime singular point. Since $T<T_{Q}$, $p\in Q^{c}$. Let $W$ be an open ball in $Q^{c}$ centered at $p$. By Theorems 13, 14, and 15, $\operatorname{genus}(M(t)\cap W)=0$ for all regular times $t<T$. Thus $(p,T)$ must be a genus 00 singularity of the flow. Thus $T<T_{\textnormal{pos}}$, since at time $T_{\textnormal{pos}}$, there must be a singularity of positive genus.
@@
Suppose that $M\in\mathscr{A}$ and that $T:=T_{Q}(M)<\infty$. Let $p\in M(T)\cap Q$ and let $\Sigma$ be a shrinker at the spacetime point $(p,T)$.
@@
If $(p,T)$ were a regular point of the flow, then for all $t<T$ sufficiently close to $T$, there would be a unique point $p(t)$ in $M(t)$ closest to $p$. But, since $p(t)\notin Q$, the image of $p(t)$ under rotation by $\pi$ about the line $\{sp:s\in\mathbf{R}\}$ is another point in $M(t)$ closest to $p$, a contradiction. Hence $(p,T)$ is a singular spacetime point of the flow.
Let $\sigma$ be the genus of the singular point $(p,T_{Q})$. Let $S$ be the set consisting of the $g+1$ points in $Q$ at distance $|p|$ from 00, and let
@@
Recall that $\Sigma$ is a smooth, properly embedded surface. If $0\in\Sigma$, then the oriented tangent plane to $\Sigma$ would be invariant under $G$. But there is no such plane. Hence $0\notin\Sigma$.
If $q$ were a point in $Q\cap\Sigma$, then (as we have just shown) $q\neq 0$. The oriented tangent plane to $\Sigma$ at $q$ is invariant under the stabilizer subgroup $G_{q}$, which implies that $\Sigma$ intersects the ray $\mathbf{R}^{+}p$ orthogonally at $p$. But that is impossible since $\Sigma$ is a smooth limit of surfaces disjoint from $Q$.
We have shown that $\Sigma\subset Q^{c}$. Also, it has genus $\leq g$. It is connected since any smooth embedded shrinker in $\mathbf{R}^{3}$ must be connected. Finally, by the convex hull property, 00 is in the convex hull of $\Sigma$. Thus $\Sigma$ has genus $g$ by Theorem 15. ∎
@@
Suppose $M_{n}$ are compact, smoothly embedded surfaces in $Q^{c}$ such that $T_{Q}(M_{n})=\infty$. Suppose the $M_{n}$ converge smoothly to a surface $M$ in $\mathscr{A}$. Then

### v3

@@ Abstract.
F. Martín’s research is partially supported by the framework of IMAG-María de Maeztu grant CEX2020-001105-M and by the grant PID2024-156031NB-I00 both funded by MCIU/AEI/ 10.13039/501100011033. Some of this work was carried out while the authors were visitors at the Simons Laufer Mathematical Sciences Institute (SLMath, formerly MSRI) in Berkeley, CA during the Fall 2024 semester, in a program supported by National Science Foundation Grant No. DMS-1928930.
@@ 1. Introduction
Figure 1. A numerical simulation of the shrinkers constructed in Section 7, for genus $g=1,\ldots,6$. Figures courtesy of Francisco Torralbo (UGR).
@@ 2. Preliminaries
If $K$ is a closed subset of $\mathbf{R}^{3}$, or, more generally, of a smooth Riemannian $3$-manifold NN, we let $K(t)$ denote the result of flowing $K$ for time $t$ by the level set flow. Let $\operatorname{T_{\textnormal{fat}}}(K)$ be the fattening time, i.e, the infimum of times $t$ such that $K(t)$ has nonempty interior.
If (as we are assuming in this section) $M$ is compact and smoothly embedded, then $\operatorname{T_{\textnormal{fat}}}(M)$ is greater than or equal to the first time when classical mean curvature starting from $M$ develops a singularity. In particular, $\operatorname{T_{\textnormal{fat}}}(M)>0$.
By [bamler-kleiner], $M(t)$ is a compact, smoothly embedded surface for almost every $t\in[0,\operatorname{T_{\textnormal{fat}}}]$, and the set of such times is relatively open in $0,\operatorname{T_{\textnormal{fat}}}]$. Furthermore,
($\star$) $t\in[0,\operatorname{T_{\textnormal{fat}}}]\mapsto\mu(t):=\mathscr{H}^{2}\llcorner M(t)$
is a unit-regular Brakke flow.
(Roughly speaking, a unit-regular Brakke flow is an integral Brakke flow that is smooth (without sudden vanishing) in a spacetime neighborhood of each point of Gauss density $1$. See [white-boundary]*Definition 11.1???? for the precise definition. Weak limits of unit-regular flows are also unit regular [white-local].)
By slight abuse of terminology, we will refer to
$t\in[0,\operatorname{T_{\textnormal{fat}}}]\mapsto M(t)$
as a Brakke flow, when we mean the Brakke flow ($\star$2).
By [bamler-kleiner], if $0\leq t_{1}<t_{2}\leq\operatorname{T_{\textnormal{fat}}}(M)$ are times at which $M(t)$ is smooth, then
@@
By [bamler-kleiner], all tangent flows at times $t\in(0,\operatorname{T_{\textnormal{fat}}}]$ are given by smoothly embedded shrinkers of multiplicity $1$. Thus if $p\in N$ and $T\in(0,\operatorname{T_{\textnormal{fat}}}(M)]$, then every sequence $t_{i}\uparrow T$ has a subsequence such that
@@
converges smoothly and with multiplicity $1$ on compact subsets to a shrinker $\Sigma$. If $\tau_{i}/t_{i}$ is bounded away from 0 and away from $\infty$, then
@@
We expect there to be only one shrinker at each spacetime point, but that is not known in all cases. (See [chodosh-schulze] for some cases in which uniqueness is known.) However, all the shrinkers at $(p,T)$ have the same genus (by Theorem 68 in the appendix), so the following definition makes sense.
@@
At a singular point of genus 0, the shrinker is either a sphere or a cylinder [brendle]. (Earlier, Kleene and Møller proved the result assuming an axis of rotational symmetry [kleene-moller].) Furthermore, the shrinker is unique by [colding-m]. (See also [zhu].)
@@
Theorem 5.
If $\operatorname{T_{\textnormal{fat}}}<\infty$, then there is shrinker at time $\operatorname{T_{\textnormal{fat}}}$ with genus $\geq 1$.
Proof.
By [chh]*Theorem 1.9, there is a shrinker $\Sigma$ at time $\operatorname{T_{\textnormal{disc}}}\leq\operatorname{T_{\textnormal{fat}}}$ that is not a plane, a shrinking sphere, or a cylinder. (See [hw]*Definition 2 or [chh]*p. 227 for the definition of the discrepancy time $\operatorname{T_{\textnormal{disc}}}=\operatorname{T_{\textnormal{disc}}}(M)$.) Since planes, spheres, and cylinders are the only smooth genus 0 shrinkers, $\Sigma$ must have genus $\geq 1$. Finally, by [bamler-kleiner]*Theorem 1.8, $\operatorname{T_{\textnormal{disc}}}=\operatorname{T_{\textnormal{fat}}}$. ∎
@@
The time $T_{\textnormal{pos}}(M)$ is the infimum of times $t\leq\operatorname{T_{\textnormal{fat}}}$ at which there is a singularity of positive genus.
In some papers, $T_{\textnormal{pos}}(M)$ is written $T_{\textnormal{gen}}(M)$.
Theorem 7.
If $T_{\textnormal{pos}}<\infty$, then there is a shrinker at time $T_{\textnormal{pos}}$ of genus $\geq 1$.
Proof.
Note that, by definition,
$T_{\textnormal{pos}}(M)\leq\operatorname{T_{\textnormal{fat}}}(M).$
If $T_{\textnormal{pos}}=\operatorname{T_{\textnormal{fat}}}$, then there is a positive-genus shrinker at time $T_{\textnormal{pos}}$ by Theorem 5. If $T_{\textnormal{pos}}<\operatorname{T_{\textnormal{fat}}}$, then there is a positive-genus shrinker at time $T_{\textnormal{pos}}$ because the set of spacetime points at which the genus is positive is a closed subset of spacetime. See, for example, [white-genus]*Theorem 25. ∎
Remark.
In this paper, we only need to flow up to time $T_{\textnormal{pos}}$.
@@
Suppose $t\in[0,T)\mapsto M(t)$ is a mean curvature flow in an open set $U$ such that $M(0)$ is a smooth, properly embedded, multiplicity $1$ submanifold of $U$. Let $K$ be a compact region in $U$ with smooth boundary transverse to $M(0)$. There is an $\epsilon>0$ with the following properties.
@@
$K^{\prime}$ $=\{\operatorname{dist}(\cdot,K)\leq\delta\},$
@@
Suppose $p_{i}$ converges to $p\in M(0)$ and that $t_{i}>0$ converges to 0. Let $q_{i}$ be the point in $M(0)$ closest to $p_{i}$. Let
@@
($\dagger$) The flow $t\in[0,2\epsilon]\mapsto M(t)\cap K^{\prime}$ is smooth.
Assertion (1) follows immediately. Assertion (2) follows from ($\dagger$2) and the Local Regularity Theorem (Theorem 8).
@@ 3. Oriented Flows
Suppose that $t\in I\mapsto M(t)$ is a unit-regular Brakke flow. We let $\operatorname{Reg}(t)$ be the set of points $p$ such that the flow is smooth and with multiplicity $1$ in a spacetime neighborhood of $(p,t)$. Equivalently, $\operatorname{Reg}(t)$ is the set of points $p$ such that the Gauss Density at $(p,t)$ is $1$. We let $\operatorname{Sing}(t)$ be the set of $p$ such that the Gauss Density of at $(p,t)$ is >1>1.
@@ 4. Platonic and Regular Polygon Shrinkers
    $\Sigma$ occurs as a shrinker in a mean curvature flow $M(\cdot)$ in $\mathbf{R}^{3}$ for which $M(0)$ is a compact, smoothly embedded surface in $Q^{c}$.
@@
If $M\in\mathscr{A}$ and if $t<\min\{T_{Q},T_{\textnormal{pos}}\}$ is a regular time of the flow $M(\cdot)$, then $M(t)\in\mathscr{A}$.
Later we will see that $T_{Q}=T_{\textnormal{pos}}$.
@@
The theorem is a trivial consequence of the facts that (1) level set flow preserves symmetries, and (2) for any compact smooth embedded surface, $\operatorname{genus}(M(t))$ is a decreasing function of regular times in $0,T_{\textnormal{pos}}]$. (See (1).) ∎
@@
If $M\in\mathscr{A}$, then there is a connected component $\Sigma$ of $M$ such that the convex hull of $\Sigma$ contains points of $Q$.
@@
Let $M$ be a compact, $G$-invariant surface embedded in $\mathbf{R}^{3}$, disjoint from $Q$, and of genus $\leq g$. Suppose $M$ has a connected component $\Sigma$ whose convex hull contains points of $Q$. Then $\Sigma$ is a $G$-invariant, genus $g$ surface, and if $\Omega$ is a simply connected open subset of $Q^{c}$, then $\operatorname{genus}(M\cap\Omega)=0$.
Proof of Theorem 17.
@@
$\Psi$ contains points in $F_{12}$ and points in $F_{13}$.
@@
Suppose not. Then (by relabelling) we may assume that $\Psi$ contains no points in the face $F_{12}$. It follows, by iterated reflections in the planes in $\mathscr{P}$ containing $E_{3}$, that $\Sigma$ lies in $\operatorname{interior}(K_{3})$, a convex open set disjoint from $Q$. But that contradicts the hypothesis that the convex hull of $\Sigma$ contains points of $Q$. Thus the claim is proved.
@@
Let $\gamma$ be an embedded curve in $\Psi$ such that one of its endpoints is in $F_{12}$ and the other endpoint is in $F_{13}$. Note that $\gamma$ is disjoint from $E_{1}$ since $M\subset Q^{c}$. Iterated reflections in the planes of $\mathscr{P}$ containing $E_{1}$ produces a simple closed curve $C$ in $\Sigma^{*}\cap\operatorname{interior}(K_{1})$ that winds once around $E_{1}$ in $K_{1}\setminus E_{1}$. In particular,
[…]
Since $C\subset\Sigma$ is homologically nontrivial in $Q^{c}$, it is homologically nontrivial in $\Sigma$, and thus
$\operatorname{genus}(\Sigma)\geq 1.$
Claim 4.
$\Psi$ contains points in $F_{23}$.
Suppose not. It follows, by iterated reflections in the planes in $\mathscr{P}$ containing $E_{1}$, that $\Sigma$ lies in $K_{1}$. For each ray $E_{i}^{\prime}$ of $Q$, $M$ contains a copy of $\Sigma$. There are $g+1$ such rays, so $M$ contains $g+1$ such copies. Since each copy has positive genus, $M$ has genus $\geq g+1$, a contradiction. Thus the claim is proved.
Since $\Sigma$ contains points in each $F_{ij}$, it follows that $\Sigma$ is invariant under reflection in each of the corresponding three planes. Those reflections generate $G$, so $\Sigma$ is invariant under $G$.
Now $E_{1}$ is one of the rays of $Q$. Let $E_{1}^{\prime}$ be another ray of $Q$, and let $\sigma\in G$ map $E_{1}$ to $E_{1}^{\prime}$. Then $C^{\prime}:=\sigma(C)$ is a simple closed curve in $\Sigma\cap\operatorname{interior}(K_{1}^{\prime})$, where $K_{1}^{\prime}=\sigma(K_{1})$, and $C^{\prime}$ winds once around $E_{1}^{\prime}$ in $K_{1}^{\prime}\setminus E_{1}^{\prime}$. Note that $C$ and $C^{\prime}$ are disjoint since $\operatorname{interior}(K_{1})$ and $\operatorname{interior}(K_{1}^{\prime})$ are disjoint.
[…]
Since $\Sigma\setminus\cup\mathscr{G}^{\prime}$ is connected,
$0\leq\operatorname{genus}(\Sigma\setminus\cup\mathscr{G}^{\prime})=\operatorname{genus}(\Sigma)-g\leq\operatorname{genus}(M)-g\leq 0.$
Thus
(3) $\operatorname{genus}(\Sigma)=\operatorname{genus}(M)=g,$
and
$\operatorname{genus}(\Sigma\setminus\mathscr{G}^{\prime})=0.$
By (3), $M\setminus\Sigma$ is a union of spheres, so
(4) $\operatorname{genus}(M\setminus\mathscr{G}^{\prime})=\operatorname{genus}(\Sigma\setminus\mathscr{G}^{\prime})=0.$
We have shown that $\Sigma$ is a $G$-invariant surface of genus $g$, and also that the genus of $M$ is equal to $g$. It remains to prove the last assertion of the theorem.
Let $U$ be the region bounded by $M$ that is disjoint from $Q$. (Thus, a point in $M^{c}$ is in $U$ if and only if there is a path from that point to a point in $Q$ that is transverse to $M$ and that intersects $M$ in an odd number of points.) Let $K=\overline{U}=U\cup M$. Thus $K$ is smooth 3-manifold-with-boundary, the boundary being $M$. Since $K$ is a smooth, compact $3$-manifold-with-boundary in $\mathbf{R}^{3}$, we can use [hatcher]*Lemma 3.5 and Poincaré duality to get that
$\dim H_{1}(K)\geq\frac{1}{2}\dim H_{1}(M)=g.$
On the other hand, one can easily get the reverse inequality using the Mayer-Vietoris Theorem. Hence, we deduce that:
$\dim H_{1}(K)=\frac{1}{2}\dim H_{1}(M)=g.$
Now consider the inclusions
$\mathscr{G}^{\prime}\subset K\subset Q^{c}.$
Since $\mathscr{G}^{\prime}$ is a basis for $H_{1}(Q^{c})$, and since $H_{1}(\mathscr{G})$, $H_{1}(K)$, and $H_{1}(Q^{c})$ each have dimension $g$, the inclusion $\iota$ of $K$ into $Q^{c}$ induces an isomorphism on $H_{1}$:
(5) $\iota_{\#}:H_{1}(K)\xrightarrow{\cong}H_{1}(Q^{c}).$
@@
It remains to show that if $\Omega$ is a simply connected open subset of $Q^{c}$, then $M\cap\Omega$ has genus 0. Suppose, to the contrary, that $M\cap\Omega$ has positive genus. Then, by elementary topology (see Lemma 19 below), $M\cap\Omega$ contains a simple closed curve $\alpha$ that does not bound a surface in $K$. Therefore $\alpha$ does not bound a surface in $Q^{c}$ since the map (5) is an isomorphism, and therefore $\alpha$ does not bound a surface in $\Omega$. But that contradicts the assumption that $\Omega$ is simply connected. ∎
The following theorem gives additional information about the surface in Theorem 17.
Theorem 18.
Let $M$ be as in Theorem 17 and let $R>0$. Suppose the convex hull of a connected component of $M\cap B(0,R)$ contains points in $Q$. Then for every $r>R$, $M\cap B(0,r)$ has genus 0 or genus $g$.
Proof.
It suffices to prove Theorem 18 in the case when $\partial B(0,r)$ intersects $M$ transversely.
Case 1: The convex hull of one of the components of $M\cap\partial B(0,r)$ contains a point of $Q$. Let $C$ be such a component, and let $\mathscr{G}$ be the union of its images under the group $G$. Exactly as in the proof of Theorem 17 (see (4)), $M\setminus\cup\mathscr{G}$ has genus 0. Thus $M\cap B(0,r)$ has genus 0.
Case 2: The convex hull of each component of $M\cap\partial B(0,r)$ is disjoint from $Q$. Then each such component bounds an embedded disk in
$\mathbf{R}^{3}\setminus(B(0,r)\cup Q).$
We can choose those disks so that the union $M^{\prime}$ of those disks and $M\cap B(0,r)$ is a smoothly embedded, $G$-invariant surface. By Theorem 17, $M^{\prime}$ has genus $g$. Thus
$\operatorname{genus}(M\cap B(0,r))=\operatorname{genus}(M^{\prime})=g.$
@@
Suppose that $K$ is a smooth, $3$-dimensional manifold-with-boundary. Let $M=\partial K$. Suppose that $W$ is a relatively open subset of $K$ such that $M\cap W$ has positive genus. Then $M\cap W$ contains a simple closed curve that is homologically nontrivial in $K$.
@@
Let $\alpha_{1}$ and $\alpha_{2}$ be a pair of simple closed curves in $M\cap W$ that intersect transversely and at exactly one point. We claim that at most one of $\alpha_{1}$ and $\alpha_{2}$ can be homologically trivial in $K$. For otherwise, there would be smooth, compact embedded surfaces $\sigma_{i}$ in $K$ such that $\partial\sigma_{i}=\alpha_{i}$. We can choose the $\sigma_{i}$ so that they intersect transversely and so that they are never tangent to $M$. Thus $\sigma_{1}\cap\sigma_{2}$ is a finite collection of curves, and so they have an even number of endpoints. But the set of their endpoints is $\alpha_{1}\cap\alpha_{2}$, which has only one point, a contradiction. ∎
@@
Suppose $M\in\mathscr{A}$. (See Definition 14.) Then $T_{Q}\leq T_{\textnormal{pos}}$.
@@
Let $0<T<\min\{T_{Q},T_{\textnormal{pos}}\}$. Let $(p,T)$ be a spacetime singular point. Since $T<T_{Q}$, $p\in Q^{c}$. Let $W$ be an open ball in $Q^{c}$ centered at $p$. By Theorems 15, 16, and 17, $\operatorname{genus}(M(t)\cap W)=0$ for all regular times $t<T$. Thus $(p,T)$ must be a genus 0 singularity of the flow. Thus $T<T_{\textnormal{pos}}$, since at time $T_{\textnormal{pos}}$, there must be a singularity of positive genus.
@@
Suppose that $M\in\mathscr{A}$, and, therefore, that $T:=T_{Q}(M)<\infty$. Let $p\in M(T)\cap Q$ and let $\Sigma$ be a shrinker at the spacetime point $(p,T)$.
@@
If $(p,T)$ were a regular point of the flow, then for all $t<T$ sufficiently close to $T$, there would be a unique point $p(t)$ in $M(t)$ closest to $p$. Now the group $G$ contains a nontrivial group of rotations about the line $\{sp:s\in\mathbf{R}\}$. Since $p(t)\notin Q$, the image of $p(t)$ under those rotations would be other points in $M(t)$ closest to $p$, a contradiction. Hence $(p,T)$ is a singular spacetime point of the flow.
Let $\sigma$ be the genus of the singular point $(p,T_{Q})$. Let $S$ be the set consisting of the $g+1$ points in $Q$ at distance $|p|$ from 0, and let
@@
First, we show that $\Sigma$ is disjoint from $Q$. Note that there exist $t_{i}\uparrow T$ such that the surfaces
$M_{i}^{\prime}:=(T-t_{i})^{-1/2}M(t_{i})$
converge smoothly and with multiplicity one to $\Sigma$. Thus if $q\in\Sigma$, then (for large $i$) there is a unique point $q_{i}$ in $M_{i}^{\prime}$ closest to 0. But if $q$ were in $Q$, then, since $M_{i}^{\prime}$ is disjoint from $Q$, the symmetries of $M_{i}^{\prime}$ imply that there are multiple points in $M_{i}^{\prime}$ closest to $q$, a contradiction. Hence $q\notin Q$. Thus $Q$ and $\Sigma$ are disjoint.
By Theorem 18, for each $i$ and each $r>0$,
$M_{i}^{\prime}\cap B(0,r)$
has genus 0 or genus $g$. Thus, by smooth convergence of $M_{i}^{\prime}$ to $\Sigma$, $\Sigma$ has genus 0 or genus $g$.
If $\Sigma$ had genus 0, then, by Brendle’s Theorem [brendle]. it would be a plane through the origin, a cylinder whose axis contains the origin, or a sphere centered at the origin, all of which intersect $Q$. But $\Sigma$ is disjoint from $Q$. Thus, $\Sigma$ has genus $g$. ∎
@@
Suppose $M_{n}$ are compact, smoothly embedded surfaces in $Q^{c}$ such that $T_{Q}(M_{n})=\infty$. Suppose the $M_{n}$ converge smoothly to a surface $M$ in $\mathscr{A}$. (See Definition 14.) Then